OpenStudy (anonymous):

he price p of a certain commodity and the quantity x sold obey the demand equation p = 600 - x/70. Suppose that the cost C of producing x units is given by C = 72000 + 60 x. Express the revenue R as a function of the number sold. (Hint: revenue = number sold * price.) R(x) =

OpenStudy (anonymous):

R = 600x-x^2/70-72000x-60x^2

OpenStudy (mertsj):

How did you get that so fast?

OpenStudy (anonymous):

no idea

OpenStudy (radar):

In other words why wouldn't R = 600x -( x^2)/70 - 72000-60x

OpenStudy (radar):

Why was the cost which is expressed in terms of x (the quantity of the commodity) multiplied by the quantity again?

OpenStudy (anonymous):

R = 600- x/70 - 72000-60x

OpenStudy (radar):

Yes cinar, I messed up and mulltiplie the price by x when in fact it too is expressed in terms of x.

OpenStudy (anonymous):

do i keep x/70 in a fraction form?

OpenStudy (radar):

What else are you considering to do with it? Do you intend to convert it to a decimal?

OpenStudy (anonymous):

it does not matter..

OpenStudy (radar):

It simpler the way it is.

OpenStudy (anonymous):

well after that i have to find the equation for price and maximize it

OpenStudy (anonymous):

thats why i asked

OpenStudy (anonymous):

hold on r(x)= 600- x/70 - 72000-60x? i think i need to get x by itself right?

OpenStudy (radar):

The price P = 600- (1/70)x or is it supposed to be 600- (1)/(70x) ?

OpenStudy (anonymous):

well im doing homework online and when i plug it in its wrong

OpenStudy (radar):

Is the x in the denominator? See my last question?

OpenStudy (anonymous):

yes

OpenStudy (radar):

Well that changes things. From the problem at the top box it appears in the numerator. Please verify.

OpenStudy (radar):

I am not sure about this problem, because when the P formula is differentiated P' I cannot solve for a minima. I would repost the problem and insure it is clearly and correctly stated. I have to go right now. good luck with it.

OpenStudy (anonymous):

ok thanks